Grade 11Physical Chemistry

Ground state energy of H-atom is (-E1), the velocity of photoelectrons emitted when photon of energy E2 strikes stationary Li2+ ion in ground state will be?

Praneet Debnath
10 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Energy conversation principle:
-E1 + ½ mv^2 = E2
1/2mv^2 = E2 + E1
v^2 = 2(E2+E1)/m
v = root(2(E2+E1)/m)